# OmniVox: Zero-Shot Emotion Recognition with Omni-LLMs

### v1

@@ 4 Experiments
We split our experiments into four subsections: Section 4.1 begins with results on each of the modalities (audio, text, text+audio). Section 4.2 focuses on prompting strategies for omni-LLMs, focusing exclusively on the audio modality. Section 3 performs an analysis on how many turns of context help for audio only experiments. Finally, Section 4.4 compares OmniVox’s results on all modalities to previous works state-of-the-art.
@@ 4.1 Multimodal Results
Audio GPT-4o-audio achieves the highest W-F1 (51.8) when context is added (c=3), outperforming Gemini Lite (51.5) and Gemini (49.9). In all cases for the audio modality besides Phi-4, we find context helps. However, with no-context conditions (c=0), performance for GPT-4o-audio and Gemini is comparable (45.8 and 45.6, respectively). Our audio-only, with context results beat all other modalities for IEMOCAP (with the exception of Phi-4)
Text Both Gemini and Gemini Lite show large gains with context added compared to no context (increase of 5.0% for Gemini, and 5.5% for Gemini-Lite) . However, our text-only results still under perform the audio results.
@@
Audio Unlike IEMOCAP, GPT-4o-audio achieves the highest W-F1 (51.3) in the no-context condition (c=0), substantially outperforming both Gemini (43.7) and Gemini Lite (44.2). Interestingly, when context is added (c=3), GPT-4o-audio’s performance decreases (to 47.7) while Gemini improves (to 48.9), making Gemini the best performer with context. For our open model Phi-4, we observe a slight decrease with added context (1.0%), and these results remain considerably below the closed models. We further explore how context hurts or helps in Section 3.
Text The text modality shows substantially higher performance than audio, unlike IEMOCAP. Gemini achieves the highest overall W-F1 score (62.8) with context (c=3), showing a modest improvement (1.6%) over the no-context condition. Gemini consistently outperforms Gemini Lite in both conditions.
Text+Audio For our multimodal text and audio experiments, we find Gemini-Lite performs better than text-only with context (1.4% increase) and audio-only with context (14% increase) and Gpt-4o-Audio performs significantly better than audio-only (14% increase). Interestingly, for Gemini, adding audio (with context), decreasing performance by 0.9%. We again find that Phi-4 shows significantly lower performance compared to closed models.
@@ 4.2 Prompting Strategies
Table 1 emphasizes that the Gold Feats. column, which uses text descriptions of gold acoustic features from Wu et al. (2024), achieves better or competitive results compared to audio-only. We aim to investigate the following question: does explicitly prompting for acoustic features help omni-LLMs for the audio-only modality? We present results for three different prompting strategies Minimal, Acoustic, and Chain-of-Thought (CoT) in Table 2.

### v2

@@ 4 Experiments
We split our experiments into four subsections: Section 4.1 begins with results on each of the modalities (audio, text, text+audio). Section 4.2 focuses on prompting strategies for omni-LLMs, exclusively using the audio modality. Section 3 performs an analysis on how many turns of context help for audio-only experiments. Finally, Section 4.4 compares OmniVox’s results on all modalities to previously reported results.
@@ 4.1 Multimodal Results
Audio-only, with context (c=3) results beat all other modalities for IEMOCAP, for three models. Phi-4 is an outlier, consistently performing worse than the other models under all conditions, and showing a different pattern for its best result.
Audio GPT-4o-audio with context (c=3) achieves the highest W-F1 (51.8) over all conditions and models, outperforming Gemini Lite (51.5) and Gemini (49.9). In all cases for the audio modality besides Phi-4, we find context helps. However, with no-context conditions (c=0), performance for GPT-4o-audio and Gemini is comparable (45.8 and 45.6, respectively).
Text Both Gemini and Gemini Lite show large gains with context added compared to no context (increase of 5.0% for Gemini, and 5.5% for Gemini-Lite) . However, our text-only results still are below the audio results.
@@
For MELD, unlike IEMOCAP, the text modality did increase the performance, with Gemini achieving the highest overall W-F1 score on Text only, and Gemini-Lite and Gpt-4o-Audio performing best on Text+Audio, all with context (c=3). Again, Phi-4 is an outlier, consistently performing worse than the other models, and again showing a different pattern for its best result.
Audio Unlike IEMOCAP, GPT-4o-audio achieves the highest W-F1 (51.3) in the no-context condition (c=0), substantially outperforming both Gemini (43.7) and Gemini Lite (44.2). Interestingly, when context is added (c=3), GPT-4o-audio’s performance decreases (to 47.7) while Gemini improves (to 48.9).
Text Gemini achieves the highest overall W-F1 score (62.8) with context (c=3), showing a modest improvement (1.6%) over the no-context condition. Gemini consistently outperforms Gemini Lite in both conditions.
Text+Audio Gemini-Lite performs better than text-only with context (1.4% increase) and audio-only with context (14% increase) and Gpt-4o-Audio performs better than audio-only by a large margin (14% increase).
@@ 4.2 Prompting Strategies
Table 1 emphasizes that the Gold Feats. column, which uses the text descriptions of gold acoustic features from Wu et al. (2024), achieves better or competitive results compared to audio-only. We aim to investigate the following question: does explicitly prompting for acoustic features help omni-LLMs for the audio-only modality? We present results for three different prompting strategies Minimal, Acoustic, and Chain-of-Thought (CoT) in Table 2.